An airplane at cruising altitude descending for a landing. The function f(t)=33,600-1500t describes its altitude t minutes after starting it’s descen. What is the airplanes altitude after 10 minutes
Answer by Boreal(15235)

f(10)=33600-1500(10)=33600-15000=18600 feet (presumed units).